In the Haber–Bosch process, ammonia is synthesized from hydrogen and nitrogen gas. The equilibrium concentrations of NH3, H2, and N2 are 0.0040 M, 0.20 M, and 0.080 M, respectively. Determine the directional shift of the reaction for each of the following situations: i. All three concentrations are 0.20 M ii. All three concentrations are 2.0 M
Answer #1

   N2(g)   + 3H2(g) ----------------> 2NH3(g)

[N2]   = 0.08M

[H2]   = 0.2M

[NH3]   = 0.004M

Kc    = [NH3]^2/[N2][H2]^3

        = (0.004)^2/(0.08)(0.2)^3   = 0.025

i.

Qc = [NH3]^2/[N2][H2]^3

     = (0.2)^2/(0.2)(0.2)^3

      = 25

Qc> Kc the equilibrium shift to left side ie reactant side

ii.

Qc = [NH3]^2/[N2][H2]^3

        = (2)^2/(2)(2)^3   = 0.25

Qc> Kc the equilibrium shift to left side ie reactant side